Solve the system by elimination or substitution.Show all steps.
-15x+3y=6
y=2+5x

Found 2 solutions by stanbon, haileytucki:
Answer by stanbon(75887)

-15x+3y=6
y=2+5x
----
Substitute for "y" and solve for "x":
---
-15x + 3(2+5x) = 6
-15x+6+15x = 6
6 = 6
-----
That means x can take any value.
The solution is the set of points (x, 2+5x)
---
It means the two equations are really the same equation.
It means the system of equations is dependent.
It means the graph of the two equations only gives
you one line.
==================
Cheers,
Stan H.

Answer by haileytucki(390)

-15x+3y=6_y=2+5x
Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is 5x+2.
-15x+3(5x+2)=6_y=5x+2
Multiply 3 by each term inside the parentheses.
-15x+15x+6=6_y=5x+2
Since -15x and 15x are like terms, subtract 15x from -15x to get 0.
0+6=6_y=5x+2
Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression.
6=6_y=5x+2
Since 6=6, the equation will always be true.
Always True_y=5x+2